Grade 11Mechanics

A body is thrown vertically upward from a point A 125m above the ground. It goes up to a maximum height of 250m above the ground and passes through A on its downward journey. the velocity of the body when it is at a height of 70m above the ground is (g=10m/s)

Prani
8 Years agoGrade 11

1 Answer

Venkat
8 Years ago
A body is thrown vertically upward from a point  A 125m above the ground. It goes up to a maximum height of 250m above the ground and passes through A on its downward journey. the velocity of the body when it is at a height of 70m above the ground is (g=10m/s)

At h=250 m, velocity will be zero. u = 0

v²-u² = 2as

u = 0, a = 10 s =(250-70)

v² = 2× 10 × (250–70)

v²= 20 ×180

v = 60 m/s